OpenStudy (anonymous):

what is 1.25^10 equal?

OpenStudy (anonymous):

= 1.25^5 * 1.25^5 = 3.051757812 * 3.051757812 = 9.313225743

OpenStudy (anonymous):

9.313225743

OpenStudy (phi):

If you don't have a calculator, type 1.25^10 = in the google search window

OpenStudy (phi):

btw, it means 1.25 times itself 10 times.